The product of digit of two digit number is 24.If 18 is added to the number digits are reversed.find the number.

Answer by ankor@dixie-net.com(22740)

The product of digit of two digit number is 24.
If 18 is added to the number of digits are reversed. find the number.
:
Let x = the 10's digit
Let y = the units
then
10x + y = the number
:
"The product of digit of two digit number is 24."
x * y = 24
:
"If 18 is added to the number, digits are reversed."
10x + y + 18 = 10y + x
10x - x + 18 = 10y - y
9x + 18 = 9y
simplify, divide by 9. results
x + 2 = y
:
Using the 1st equation, replace y with (x+2)
x(x+2) = 24
x^2 + 2x - 24 = 0; our old friend, the quadratic equation!
Factors to
(x+6)(x-4) = 0
the positive solution
x = 4
then
y = 4 + 2
y = 6
:
46 is the number
:
:
Check this: 46 + 18 = 64, the number reversed
:
The solution, x=-6 will also work, check it out.
-64 is the number,